OpenStudy (eva12):

1)find the partial derivatives

OpenStudy (eva12):

\[z=x^3+2^y+x^y \] the answer is\[2^yln(2)+x^yln(x)\] how do they get the answer please show your work

OpenStudy (eva12):

its zy

OpenStudy (zzr0ck3r):

\(z=x^3+2^y+x^y\implies \frac{dz}{dy}=\frac{d(x^3)}{dy}+\frac{d(2^y)}{dy}+\frac{d(x^y)}{dy}\). the derivative of a^y is ln(a)*y, so \(\frac{dz}{dy}=0+\ln(2)*2^y+\ln(x)x^y\)

OpenStudy (zzr0ck3r):

we treat x as a constant.

OpenStudy (zzr0ck3r):

that should say the derivative of a^y is ln(a)*a^y

OpenStudy (eva12):

can you please show it step by step

OpenStudy (isbella):

\[d/dy(x^3+2^y+x^y) \] \[d/dy(x^y)+d/dy(2^y)\] \[\ln(2)*2^y+\ln(x)*x^y\] \[2^yln(2)+x^yln(x)\] heres the steps hope it helps=]

OpenStudy (eva12):

thanks

OpenStudy (zzr0ck3r):

lol how is that any different?

OpenStudy (eva12):

i need the steps for my homework